If the sides of a square are lengthened by 8 cm, the area becomes 144 cm^2. Find the length of a side of the original square.
What is the length of a side of the original square is ? cm.

Answer by mananth(16946)

LET ORIGINAL LENGTH BE X
INCREASED BY 8
NEW LENGTH = X+8
Area = L^2
(x+8)^2= 144
x^+16x+64=144
x^2+16x-80=0
x^2+20x-4x-80=0
x(x+20)-4(x+20)=0
(x+20)(x-4)=0
x = 4 the original length of square
...
m.ananth@hotmail.ca